OpenStudy (anonymous):

need help solving this quadratic equation (2x-1)(x+7)=0

OpenStudy (dtan5457):

Since it is already factored, just solve for x. 2x-0 x+7=0

OpenStudy (dtan5457):

*2x-1=0

OpenStudy (acxbox22):

it is already factored for you so use the zero product theorum and do what @dtan5457 said

OpenStudy (anonymous):

@dtan5457 is there more then 1 answer to the question

OpenStudy (acxbox22):

yes there are 2 solutions

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

so for the x+7=0 i got -7 is that right